# A Minimal Subset Approach for Efficient and Scalable Loop Closure

### v1

@@ Abstract
Loop closure detection in large-scale and long-term missions can be computationally demanding due to the need to identify, verify, and process numerous candidate pairs to establish edge connections for the pose graph optimization. Keyframe sampling mitigates this by reducing the number of frames stored and processed in the back-end system. In this article, we address the gap in optimized keyframe sampling for the combined problem of pose graph optimization and loop closure detection. Our Minimal Subset Approach (MSA) employs an optimization strategy with two key factors, redundancy minimization and information preservation, within a sliding window framework to efficiently reduce redundant keyframes, while preserving essential information. This method delivers comparable performance to baseline approaches, while enhancing scalability and reducing computational overhead. Finally, we evaluate MSA on relevant publicly available datasets, showcasing that it consistently performs across a wide range of environments, without requiring any manual parameter tuning.
Index Terms:
Loop Closure, Place Recognition, Localization, Keyframe Sampling, Optimization.
@@ I Introduction
Simultaneous Localization and Mapping (SLAM) in large-scale and long-term missions presents significant challenges, particularly in maintaining accuracy and efficiency as the environment grows over time. A key component of SLAM is Pose Graph Optimization (PGO), which ensures consistent mapping by optimizing the global trajectory. However, this process relies heavily on loop closure detection, to reduce accumulated drift, which is computationally expensive, especially in large-scale environments, due to the vast number of candidate pairs that need to be identified, verified, and processed to establish node connections in the pose graph. As the mission progresses, the number of loop closure candidates grows rapidly, leading to significant delays in processing times, with some datasets generating over 150,000 candidates, requiring four and a half hours to process, exceeding that of the mission time [1]. This complexity is further compounded by the computational intensity of data association and sequence matching during map optimization, where the processing time can range from seconds to several minutes, depending on the size of the data [2, 3].
Due to these computational challenges, many SLAM systems in large-scale environments have opted to either avoid or minimize loop closure detection to reduce the burden on limited computational resources. Some approaches, particularly in expansive environments, forego loop closures entirely, as the growing number of candidates becomes increasingly difficult to handle in real-time applications, as highlighted recently in [4]. While loop closure detection is critical for correcting drift and ensuring map consistency, the computational complexity can overwhelm systems if not managed effectively, often leading to false positives that degrade the overall quality of the map [5].
@@ I-A Related Work
Efficient implementations of pose graph optimization, such as GTSAM [6], g2o [7] and iSAM2 [8] have significantly enhanced its scalability by leveraging factor graphs and incremental optimization techniques. Additionally, sparsification methods [9, 10, 11] and hierarchical approaches [12, 13] have further improved computational efficiency by reducing the number of active nodes in the optimization process. However, despite these advances, loop closure detection remains a major bottleneck due to the large number of candidate pairs that must be checked and processed to add meaningful constraints to the pose graph.
Figure 1: An example on KITTI 06. Comparison of two sampled trajectories: one using an entropy-based approach and the other using the proposed Minimal Subset Approach (MSA). The MSA-sampled trajectory achieves a lower Relative Pose Error (RPE) after the pose graph optimization.
Keyframe sampling helps reduce the computational demands of loop closure detection by limiting the number of frames stored and processed in SLAM systems. In LiDAR-(Inertial) Odometry, systems such as LIO-SAM [14, 15, 16] or back-end global optimization systems like LAMP [17, 18], use fixed Euclidean distance intervals (1–2 meters) for keyframe generation to maintain local maps and optimize global poses. Although effective, these methods lack adaptability in dynamic environments with varying densities of loop closure candidates. Recent advancements have introduced adaptive keyframe sampling techniques, adjusting intervals based on environmental spaciousness [19, 20], although manual threshold adjustment is still required. Entropy-based methods, like the one proposed by [21], use information theory to select keyframes, but face similar adaptability issues. Other approaches, such as [22], focus on displacement vector similarity to balance computational cost and map completeness while managing keyframes in a sliding window.
In this article, we address the gap in optimized keyframe sampling for pose graph optimization and loop closure detection by streamlining keyframe extraction. Our Minimal Subset Approach uses two key criteria, redundancy minimization and information preservation, within a sliding window framework to efficiently reduce redundant keyframes while maintaining essential information. This method enhances adaptability across various environments and eliminates the need for manual threshold tuning, with an example of its performance illustrated on Fig. 1.
@@ I-B Contributions
Based on the aforementioned, the contributions of this article can be summarized as follows: (a) We propose a Minimal Subset Approach (MSA) that efficiently addresses the problem of loop closure detection and pose graph optimization by leveraging a subset of keyframes. This method achieves comparable results to baseline approaches but with improved scalability and efficiency, reducing computational demands. (b) The proposed approach is environment agnostic, avoiding manual parameter tuning and offering consistent performance across diverse environments. This is achieved through a sliding window optimization in the hyper-dimensional descriptor space, eliminating the need for threshold adjustments. (c) The MSA addresses the combined problem of pose graph optimization and loop closure detection. Unlike many existing methods that focus solely on extracting keyframes for scan matching and pose graph optimization, MSA takes into account how sampling intervals impact loop closure detection performance through place recognition, providing a more integrated solution to these interrelated challenges.
@@ II-D Reducing Search Space in Place Recognition
The detection and verification of loop closure edges $\mathcal{E}_{l}\triangleq\mathcal{L}$ is computationally demanding due to the large number of potential keyframe pairs, which scales quadratically as $O(n^{2})$, where $n=|\mathbf{K}|$ is the number of keyframes. While using descriptors and methods like k-nearest neighbor (k-NN) search [26] can accelerate the process, additional steps are still required to remove outliers, mitigate false positives, or prioritize certain nodes [1], while extracting the relative pose transformation between loop closure nodes also involves computationally intensive techniques such as General Iterative Closest Point (GICP) [27] or Sample Consensus Initial Alignment (SAC-IA) [28]. This pipeline does not scale well for extensive missions, often resulting in processing times of 3–4 hours or more [4, 17, 18], and the challenge becomes even greater in multi-robot systems [2, 3].
To address this computational complexity, we propose a sampling strategy to reduce the number of keyframes considered for matching, while maintaining minimal impact on the loop closure set $\mathcal{L}$. Let $\mathbf{K}_{\text{S}}\subset\mathbf{K}$ represent a subset of keyframes selected through a sampling process $S:\mathbf{K}\rightarrow\mathbf{K}_{\text{S}}$ such that $|\mathbf{K}_{\text{S}}|\ll|\mathbf{K}|$. Our objective is to reduce the search space by confining the loop closure detection process to only the selected keyframes in $\mathbf{K}_{\text{S}}$. The goal is to identify a reduced set $\mathbf{K}_{\text{S}}^{*}$ that minimizes redundant keyframes while preserving the set of loop closures $\mathcal{L}$, a process denoted as:
@@ II-E3 Dynamic Keyframe Selection
To minimize the keyframe set size in real-time, we must consider the entire set $\mathbf{K}$; however, this is impractical for maintaining computational feasibility in loop closure detection. The challenge is to determine the contribution of each keyframe, which requires anticipating future keyframes and introduces a lack of casualty. Moreover, finding the optimal subset $\mathbf{K}_{\text{S}}^{*}$ is an NP-hard problem, involving an exhaustive search of all keyframe combinations, which becomes computationally infeasible for sets larger than 15–20 keyframes.
@@ III Minimal Subset Approach
To address these challenges, we propose a Minimal Subset Approach (MSA) to approximate a solution to the problems in Eq. (5)-(6). The proposed MSA identifies and eliminates redundancy within a keyframe set while preserving crucial information. Designed for real-time implementation, it employs a sliding window combinatorial optimization with two distinct objectives, balancing information preservation and redundancy minimization.
@@ III-A Redundancy in Keyframes
The frequency of a sensor, the platform’s speed, and the environmental characteristics can cause keyframe samples to capture redundant information if they are too closely spaced, with more details available on [30]. To address this, we define redundancy within a keyframe set and propose a metric to quantify it based on the descriptor space. A keyframe $\mathbf{k}$ is redundant within a set $\mathbf{K}$ if its removal does not affect the loop closure edge set $\mathcal{L}$ and does not create discontinuities in the map representation:
@@
$\rho_{\tau}(\mathbf{K})=\frac{1}{N-1}\sum_{i=1}^{N-1}f_{\sigma}(\mathbf{k}_{i},\mathbf{k}_{i+1}),$ (9)
@@ III-B Information Preservation in Keyframes
Let $F:\mathbb{R}^{3}\rightarrow\mathbb{R}^{M}$, represent the descriptor extraction function, either learning-based [24, 31] or handcrafted [25, 32], which maps each observation from 3D space to an $M$-dimensional representation. This function depends on the input observation (e.g. LiDAR scan), but it is understood that $\mathbf{z}_{t}\propto f(\mathbf{x}_{t})$ implies that each observation depends on the viewpoint and pose. To understand how the descriptors are sensitive to pose changes, we compute the Jacobian $\mathbf{J}$ of $F$ with respect to poses $\mathbf{x}$. Explicitly deriving the Jacobian can be impractical because of the intricate nature of functions like deep neural networks [30]. Consequently, we employ numerical approximations to estimate the rate of change, as shown in Fig. 3. Moreover, since poses are encompassed within $SE(3)$, we reduce the dimensional complexity by using the Euclidean norm for distance measurement instead of individual axis derivatives. It is essential to use yaw-invariant descriptors, as suggested in [24, 25, 33], to ensure the Jacobian calculations are meaningful by preventing orientation changes. Considering the descriptors $\mathbf{d}\in\mathbb{R}^{M}$ as random variables and poses $\mathbf{x}\in SE(3)$ as samples, the product $\mathbf{J}_{\mathbf{F}}^{\mathsf{T}}\mathbf{J}_{\mathbf{F}}$ estimates the covariance matrix:
@@
$\pi_{\tau}(\mathbf{K})=-\frac{1}{N-1}\sum_{i=1}^{N-1}f_{\delta}(\mathbf{d}_{i}^{\prime},\mathbf{d}_{i+1}^{\prime}),$ (11)
@@
Figure 3: Descriptors across poses. An example showing the descriptors, the Jacobian, and the transformed descriptors for two window keyframe sets. The Jacobian indicates how the features of the descriptors change between keyframes, while the transformed descriptors demonstrate how these same features appear after transformation to the principal components of that set.
@@ IV Experimental Setup and Results
TABLE I: Translational (t) and Rotational (rot) Absolute Trajectory Error Improvement after PGO and the False Positive Rate (FPR)
Datasets KITTI Odometry (Seq. 00 / 06 ) MulRan (DCC / KAIST) Apollo-SouthBay (SanJose / Columbia)
Metrics t. ATE [%] rot. ATE [%] FPR [%] t. ATE [%] rot. ATE [%] FPR [%] t. ATE [%] rot. ATE [%] FPR [%]
All Samples 06.9 /-78.6 -06.1 /-51.1 01.4 /-05.6 29.8 / 67.4 24.9 / 59.5 08.0 / 21.2 96.2 / 98.1 95.3 / 97.5 30.7 / 07.5
Const. 1m 12.2 /-78.7 -12.3 /-55.3 02.8 /-07.0 59.3 / 87.1 56.4 / 71.0 20.2 / 23.8 85.8 / 97.1 84.1 / 97.2 25.1 / 04.1
Const. 2m 10.5 /-44.9 -06.5 /-23.7 01.3 /-03.1 70.0 / 80.9 68.4 / 73.7 07.6 / 11.6 06.8 / 75.9 06.8 / 77.2 25.3 / 02.3
Const. 3m 05.1 /-22.6 -03.9 /-13.8 03.0 /-09.5 63.6 / 87.7 64.6 / 79.1 04.4 / 07.7 00.0 / 83.6 00.0 / 84.4 16.6 / 02.4
Entropy 09.6 /-24.1 -03.8 /-09.1 01.5 /-05.4 69.2 / 68.7 69.5 / 63.1 16.4 / 14.6 77.6 / 69.4 76.2 / 72.6 21.0 / 03.1
Spacious. 04.4 /-35.5 -02.0 /-39.1 02.3 /-06.9 68.4 / 68.6 68.1 / 63.1 08.5 / 16.1 03.5 / 67.9 03.5 / 71.7 24.1 / 04.9
MSA (Ours) 11.1 /-78.8 -09.7 /-55.5 01.1 /-04.9 70.1 / 91.4 69.7 / 77.8 08.2 / 11.4 88.7 / 84.0 88.3 / 85.1 21.5 / 01.6
We evaluate our approach using three datasets: KITTI Odometry [34], MulRan [35], and Apollo-SouthBay [36], which cover diverse environments such as urban areas, rural landscapes, and complex structures like bridges and tunnels. Each dataset provides 3D LiDAR scans and GNSS-based ground truth poses, with MulRan being particularly challenging due to obstructed LiDAR views. For descriptor extraction, we use OverlapTransformer (OT) [24] with a $1\times 256$ feature vector. In [30], we demonstrate the generalizability of our method to other descriptors, like Scan Context [25]. Comparisons are made with constant intervals of 1, 2, and 3 meters [14, 17, 18, 15, 16], as well as with adaptive sampling based on LiDAR spaciousness [19] and entropy [21]. All experiments were conducted on a 14th Gen Intel Core i9-14900K with 128GB DDR5 RAM. The parameters of MSA are set to the defaults of $\alpha=\beta=1$ and $N=10$. The window optimization is solved, on average, in 14.4ms, making it suitable for real-time sampling.

### v2

# Balancing Accuracy and Efficiency for Large-Scale SLAM: A Minimal Subset Approach for Scalable Loop Closures

@@ Abstract
Typical LiDAR SLAM architectures feature a front-end for odometry estimation and a back-end for refining and optimizing the trajectory and map, commonly through loop closures. However, loop closure detection in large-scale missions presents significant computational challenges due to the need to identify, verify, and process numerous candidate pairs for pose graph optimization. Keyframe sampling bridges the front-end and back-end by selecting frames for storing and processing during global optimization. This article proposes an online keyframe sampling approach that constructs the pose graph using the most impactful keyframes for loop closure. We introduce the Minimal Subset Approach (MSA), which optimizes two key objectives: redundancy minimization and information preservation, implemented within a sliding window framework. By operating in the feature space rather than 3-D space, MSA efficiently reduces redundant keyframes while retaining essential information. In sum, evaluations on diverse public datasets show that the proposed approach outperforms naive methods in reducing false positive rates in place recognition, while delivering superior ATE and RPE in metric localization, without the need for manual parameter tuning. Additionally, MSA demonstrates efficiency and scalability by reducing memory usage and computational overhead during loop closure detection and pose graph optimization.
@@ I Introduction
Simultaneous Localization and Mapping (SLAM) in large-scale missions presents significant challenges, particularly in maintaining accuracy and efficiency as the environment grows over time. A key component of back-end SLAM is Pose Graph Optimization (PGO), which ensures consistent mapping by optimizing the global trajectory. However, this process relies heavily on loop closure detection, to reduce accumulated drift, which is computationally expensive, especially in large-scale environments, due to the vast number of candidate pairs that need to be identified, verified, and processed to establish node connections in the pose graph. As the mission progresses, the number of loop closure candidates grows rapidly, leading to significant delays in processing times, with some datasets generating over 150,000 candidates, requiring four and a half hours to process, exceeding that of the mission time [1]. This complexity is further compounded by the computational intensity of data association and sequence matching during map optimization, where the processing time can range from seconds to several minutes, depending on the size of the data [2, 3]. Due to these computational challenges, many SLAM systems in large-scale environments have opted to either avoid or minimize loop closure detection to reduce the burden on limited computational resources. Some approaches, particularly in expansive environments, forego loop closures entirely, as the growing number of candidates becomes increasingly difficult to handle in real-time applications, as highlighted recently in [4]. While loop closure detection is critical for correcting drift and ensuring map consistency, the computational complexity can overwhelm systems if not managed effectively, often leading to false positives that degrade the overall quality of the map [5].
Figure 1: An example on KITTI 06. Comparison of two trajectories from KISS-ICP, after Pose Graph Optimization (PGO). In the top figure, keyframes are sampled using an entropy-based approach, while in the bottom figure, keyframes are sampled with the proposed Minimal Subset Approach (MSA) [6] which achieves a lower Relative Pose Error (RPE).
@@ I-A Related Work
Efficient implementations of pose graph optimization have significantly enhanced its scalability by leveraging factor graphs and incremental optimization techniques. In more detail, GTSAM [7] models probabilistic relationships using factor graphs and solves large-scale problems efficiently with solvers like iSAM2 [8], which incrementally updates the SLAM solution in real-time through a Bayes tree representation. Similarly, g2o [9] provides a versatile graph optimization framework, leveraging sparse pose graph connectivity for efficiency while supporting various parameterizations and loss functions. Furthermore, sparsification methods [10, 11, 12] and hierarchical approaches [13, 14] have further improved efficiency by reducing the number of active nodes in the optimization process. However, despite these advances, loop closure detection remains a major bottleneck due to the large number of candidate pairs that must be checked and processed to add meaningful constraints to the pose graph.
Keyframe sampling helps reduce the computational demands of loop closure detection by limiting the number of frames stored and processed in SLAM pipelines. In LiDAR-(Inertial) Odometry, systems such as LIO-SAM [15] and DiSCo-SLAM [16] or back-end global optimization systems like LAMP [17], use fixed Euclidean distance intervals (1–2 meters) for keyframe generation to maintain local maps and optimize global poses. Although effective, these methods lack adaptability in dynamic environments with varying densities of loop closure candidates. Recent advancements have introduced adaptive keyframe sampling techniques, adjusting intervals based on environmental spaciousness [18, 19], although manual threshold adjustment is still required. Entropy-based methods, like the one proposed by Zeng et al. [20], use information theory to select keyframes, but face similar adaptability issues. Other approaches, such as the one proposed by Ou et al. [21], focus on displacement vector similarity to balance computational cost and map completeness while managing keyframes in a sliding window.
In contrast to the above that focus solely on back-end optimization, reducing the number of nodes in pose graph optimization, or generating keyframes for submap creation, our work centers on sampling keyframes online to reduce memory allocation and loop closure query times while preserving retrieval performance during place recognition and maintaining localization accuracy after pose graph optimization. In this work, we evaluate the impact of our previously introduced optimized keyframe sampling method [6] on critical back-end tasks, including loop closure detection and pose graph optimization, with a glimpse of the results illustrated in Fig. 1. The proposed Minimal Subset Approach (MSA) employs two key criteria, redundancy minimization and information preservation, within a sliding window optimization framework to streamline keyframe selection. By reducing redundant information while preserving essential data, MSA enhances scalability and computational efficiency for both place recognition and pose graph optimization. Unlike traditional methods that rely on spatial heuristics and 3-D space, MSA utilizes the feature space, making it inherently environment-agnostic. Its reliance on the descriptor extraction framework used for loop closure detection eliminates the need for manual parameter tuning, ensuring consistent performance across diverse scenarios.
@@ I-B Contributions
Based on the aforementioned, the main contributions of this article can be summarized as follows: (a) To the best of our knowledge, this is the first paper to present the pose graph optimization (PGO) problem with an emphasis on the interplay between place recognition, keyframe sampling and loop closure detection. We highlight the importance of addressing these interconnected problems jointly to achieve a more comprehensive solution and provide further insights on the inherent challenges. (b) This paper significantly extends our previous work [6] by demonstrating the impact of keyframe sampling not only in place recognition but on pose graph optimization and metric localization as well. This provides valuable insights into the interconnection between the front-end and back-end SLAM, showcasing how sampling intervals influence the performance of the latter. (c) Through our results we demonstrate that MSA enhances scalability and efficiency by reducing computational demands. It consistently achieves lower false positive rates in place recognition and superior ATE and RPE in metric localization across diverse environments, without manual parameter tuning. By minimizing loop closure query times, memory usage, and the overall time required for pose graph optimization, MSA is well-suited for large-scale missions.
@@ II-D Reducing Search Space in Place Recognition
The detection and verification of loop closure edges $\mathcal{E}_{l}\triangleq\mathcal{L}$ is computationally demanding due to the large number of potential keyframe pairs, which scales quadratically as $O(n^{2})$, where $n=|\mathbf{K}|$ is the number of keyframes. While using descriptors and methods like k-nearest neighbor (k-NN) search [25] can accelerate the process, additional steps are still required to remove outliers, mitigate false positives, or prioritize certain nodes [1], while extracting the relative pose transformation between loop closure nodes also involves computationally intensive techniques such as General Iterative Closest Point (GICP) [26]. This pipeline does not scale well for extensive missions, often resulting in processing times of 3-4 hours or more [4, 17], and the challenge becomes even greater in multi-robot systems [3].
To address this computational complexity, we propose a sampling strategy to reduce the number of keyframes considered for matching, while maintaining minimal impact on the loop closure set $\mathcal{L}$. Let $\mathbf{K}_{\text{S}}\subset\mathbf{K}$ represent a subset of keyframes selected through a sampling process $S:\mathbf{K}\rightarrow\mathbf{K}_{\text{S}}$ such that $|\mathbf{K}_{\text{S}}|<|\mathbf{K}|$. Our objective is to reduce the search space by confining the loop closure detection process to only the selected keyframes in $\mathbf{K}_{\text{S}}$. The goal is to identify a reduced set $\mathbf{K}_{\text{S}}^{*}$ that minimizes redundant keyframes while preserving the set of loop closures $\mathcal{L}$, a process denoted as:
@@ II-E3 Dynamic Keyframe Selection
To minimize the keyframe set size in real-time, we must consider the entire set $\mathbf{K}$; however, this is impractical for maintaining computational feasibility in loop closure detection. The challenge is to determine the contribution of each keyframe, which requires anticipating future keyframes and introduces a lack of causality. Moreover, finding the optimal subset $\mathbf{K}_{\text{S}}^{*}$ is an NP-hard problem, involving an exhaustive search of all keyframe combinations, which becomes computationally infeasible for sets larger than 15–20 keyframes.
@@ III Minimal Subset Approach
To address these challenges, we propose the Minimal Subset Approach (MSA) to approximate a solution to the problems in Eq. (5)-(6). Unlike our previous work [6], which focuses solely on place recognition retrieval performance, in this article we extend the analysis to assess its impact on the subsequent step described in Eq.(7) and its comparison to the original Eq. (2). The proposed MSA identifies and eliminates redundancy within a keyframe set while preserving essential data. Designed for real-time use, it employs a sliding window combinatorial optimization with dual objectives: balancing information preservation and redundancy minimization.
@@ III-A Redundancy in Keyframes
The frequency of a sensor, the platform’s speed, and the environmental characteristics can cause keyframe samples to capture redundant information if they are too closely spaced. To address this, we define redundancy within a keyframe set and propose a metric to quantify it based on the descriptor space. A keyframe $\mathbf{k}$ is redundant within a set $\mathbf{K}$ if its removal does not affect the loop closure edge set $\mathcal{L}$ and does not create discontinuities in the map representation:
@@
$\rho_{\tau}(\mathbf{K})=\frac{1}{N}\sum_{i=0}^{N-1}f_{\sigma}(\mathbf{k}_{i},\mathbf{k}_{i+1}),$ (9)
@@ III-B Information Preservation in Keyframes
Let $F:\mathbb{R}^{3}\rightarrow\mathbb{R}^{M}$, represent the descriptor extraction function, either learning-based [23, 28] or handcrafted [24, 29], which maps each observation from 3-D space to an $M$-dimensional representation. This function does not directly depend on the pose $\mathbf{x}_{t}$, but rather on the input observation $\mathbf{z}_{t}$ (e.g. LiDAR scan), which itself depends on the pose through the sensor model $\mathbf{z}_{t}\propto h(\mathbf{x}_{t},\mathbf{W})$, where $\mathbf{W}$ represents the world. To understand how the descriptors are sensitive to pose changes, we can compute the Jacobian $\mathbf{J}$ of $F$ with respect to poses $\mathbf{x}$ through the chain rule $(\partial\mathbf{F}\,\partial\mathbf{z})/(\partial\mathbf{z}\,\partial\mathbf{x})$. Explicitly deriving the Jacobian can be impractical because of the intricate nature of functions like deep neural networks. Consequently, we employ numerical approximations to estimate the rate of change. Moreover, since poses are encompassed within $SE(3)$, we reduce the dimensional complexity by using the Euclidean norm for distance measurement instead of individual axis derivatives. It is essential to use yaw-invariant descriptors, as suggested in [23, 24], to ensure the Jacobian calculations are meaningful by preventing orientation changes. For a comprehensive discussion of these assumptions and derivations, we refer the reader to our previous work [6], which provides detailed explanations. Considering the descriptors $\mathbf{d}\in\mathbb{R}^{M}$ as random variables and poses $\mathbf{x}\in SE(3)$ as samples, the product $\mathbf{J}_{\mathbf{F}}^{\mathsf{T}}\mathbf{J}_{\mathbf{F}}$ estimates the covariance matrix:
@@
$\pi_{\tau}(\mathbf{K})=-\frac{1}{N}\sum_{i=0}^{N-1}f_{\delta}(\mathbf{d}_{i}^{\prime},\mathbf{d}_{i+1}^{\prime}),$ (11)
@@ IV Experimental Setup and Results
TABLE I: Translational (t) and Rotational (R) Absolute Trajectory Error Improvement after PGO and the False Positive Rate (FPR)
Datasets KITTI Odometry (Seq. 05 / 06 ) MulRan (DCC / KAIST) Apollo-SB (SanJose / Columbia)
Metrics t. ATE [%] R. ATE [%] FPR [%] t. ATE [%] R. ATE [%] FPR [%] t. ATE [%] R. ATE [%] FPR [%]
ScanContext All Samples 72.7 / 76.1 94.4 / 49.0 06.6 / 06.4 30.2 / 69.4 24.1 / 56.3 08.8 / 20.8 92.0 / 93.7 93.7 / 92.3 29.4 / 06.8
Const. 1.0m 90.5 / 76.2 95.9 / 52.6 04.3 / 04.1 63.2 / 88.6 55.2 / 70.7 11.6 / 24.0 84.3 / 92.5 87.4 / 92.0 26.1 / 05.5
Const. 2.0m 86.2 / 62.3 94.0 / 46.4 06.0 / 05.9 72.1 / 83.3 69.4 / 72.0 12.1 / 13.3 66.7 / 79.2 59.5 / 78.6 25.9 / 03.3
Const. 3.0m 79.8 / 60.1 77.4 / 42.3 10.5 / 09.9 58.7 / 79.5 60.6 / 69.2 19.6 / 11.7 54.4 / 77.1 40.9 / 69.8 24.0 / 03.1
Entropy-based 66.9 / 46.7 93.4 / 48.3 09.1 / 05.5 68.7 / 71.8 69.0 / 66.8 14.2 / 15.6 90.1 / 80.3 89.8 / 82.9 24.5 / 03.2
Spaciousness 77.8 / 55.6 87.6 / 46.7 09.3 / 05.9 67.9 / 70.7 68.5 / 66.1 10.0 / 16.5 79.7 / 79.6 84.2 / 82.1 25.2 / 06.0
MSA (Ours) 90.3 / 76.4 96.1 / 53.0 03.3 / 03.4 71.9 / 91.9 70.1 / 76.6 06.8 / 12.3 89.9 / 87.5 91.4 / 80.3 23.3 / 02.7
OverlapTrans All Samples 75.7 / 78.6 95.9 / 51.1 05.1 / 05.6 29.8 / 67.4 24.9 / 59.5 08.0 / 21.2 96.2 / 98.1 95.3 / 97.5 30.7 / 07.5
Const. 1.0m 94.6 / 78.7 97.3 / 55.3 06.2 / 07.0 59.3 / 87.1 56.4 / 71.0 20.2 / 23.8 85.8 / 97.1 84.1 / 97.2 25.1 / 04.1
Const. 2.0m 84.4 / 44.9 97.2 / 23.7 04.3 / 03.1 70.0 / 80.9 68.4 / 73.7 07.6 / 11.6 06.8 / 75.9 06.8 / 77.2 25.3 / 02.3
Const. 3.0m 90.5 / 22.6 87.9 / 13.8 06.0 / 09.5 63.6 / 87.7 64.6 / 79.1 04.4 / 07.7 00.0 / 83.6 00.0 / 84.4 16.6 / 02.4
Entropy-based 67.9 / 24.1 93.0 / 09.1 06.1 / 05.4 69.2 / 68.7 69.5 / 63.1 16.4 / 14.6 77.6 / 69.4 76.2 / 72.6 21.0 / 03.1
Spaciousness 85.8 / 35.5 97.2 / 39.1 06.8 / 06.9 68.4 / 68.6 68.1 / 63.1 08.5 / 16.1 03.5 / 67.9 03.5 / 71.7 24.1 / 04.9
MSA (Ours) 93.2 / 78.8 97.5 / 55.5 04.0 / 04.9 70.1 / 91.4 69.7 / 77.8 08.2 / 11.4 88.7 / 84.0 88.3 / 85.1 21.5 / 01.6
The underlined results correspond to the best performing constant sampling interval, while the bold results correspond to the best adaptive sampling method.
To support the claims of our second contribution, on how the sampling interval affects performance, we conducted a series of experiments with the following setup. Descriptor extraction was performed using two approaches: the learning-based OverlapTransformer (OT[23], which generates a $1\times 256$ feature vector, and the hand-crafted Scan Context (SC[24], which produces a $20\times 60$ feature representation. These choices demonstrate the descriptor-agnostic nature of our method and its compatibility with both learned and hand-crafted descriptors. We benchmark our approach against constant sampling intervals of 1, 2, and 3 meters [15, 17, 16, 30], as well as adaptive sampling strategies based on LiDAR spaciousness [18] and entropy [20]. All experiments were conducted on a 14th Gen Intel Core i9 system with 128GB of RAM, where the sampling window optimization ($N=10$) is solved in approximately 5–15 milliseconds, making it suitable for real-time sampling.